OpenStudy (anonymous):

Simplify the expression 1 _______ 8 -3

OpenStudy (anonymous):

sorry the one goes over the line

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

Question is not clear.

OpenStudy (anonymous):

ok the 1 is on top of the line and the 8-3 is under the line this is multiple choice and the answers they give are -1/512,-1/24,24&512

OpenStudy (anonymous):

hey no one can help

OpenStudy (anonymous):

can any one please help me with this answer

OpenStudy (anonymous):

Simplify. Express the product as a radical expression 11 radical x 5 . radical x

jhonyy9 (jhonyy9):

1/8 - 1/3 = (3-8)/24 = -5/24

OpenStudy (anonymous):

mom type it out the way i taught u. ok?

jhonyy9 (jhonyy9):

or if you thought 1/(8-3) =1/5